The length of a rectangle is 3 more than its width. The area of the rectangle is 40 square centimeters. What is the length and width of the rectangle?
Answer by anantha(86)

The length of a rectangle is 3 more than its width. The area of the rectangle is 40 square centimeters. What is the length and width of the rectangle?
sol:
let the width be x
length=3 more than the width
=3+x
area of the rectangle=length*width
=(3+x)*x
=3x+x^2
according to the problem area of the rectangle=40 square centimeters
3x+x^2=40
x^2+3x-40=0
on multiplying the two numbers we get 40 and adding or subtracting we get 3
those numbers are 5 and 8
x^2+8x-5x-40=0
x(x+8)-5(x+8)=0
(x-5)(x+8)=0
x-5=0
x=5
x+8=0
x=-8
here we are taking only positive values
x=5
width=5 centimeters
length=3+x
=3+5=8 centimeters